Suppose {an}∞n=1 be a sequence of positive real numbers and 0 < x < 1. If an+1 < x·an for every n ∈ N, prove that limn→∞an = 0


1
Expert's answer
2021-02-03T14:01:24-0500

Given that "a_{n+1} < x. a_n" and "0<x<1" we have that "a_{n+1} < a_{n} \forall n \in \mathbb{N}" So as "(a_n)_{n=1}^{n= \infty}" is a sequence of positive real numbers; we have that as n gets larger and larger "a_{n}" goes to zero i.e. "\lim \limits_{n \to \infty} a_n = 0"